OpenStudy (anonymous):

This is algebra and can someone please help me on how to do this? Clueless. okay... 27(a^6)(b^4)+12(a^2)(B^3)

OpenStudy (anonymous):

\(27(a^6)(b^4)+12(a^2)(b^3) \\3a^2b^3(9a^4b+4) \)

OpenStudy (anonymous):

but how though? I mean I trust your answer it's just they did a terrible job of teaching me?

OpenStudy (anonymous):

alright, let me think about how to explain this. so when you divide exponents with the same base it works like this: \(\dfrac{a^M}{a^N}=a^{(M-N)}\) and so when we factor, we're dividing by the greatest common factor. for example the factors of 27: 1 3 9 27 12 1 2 3 4 6 12 so we divide each term by 3 and we're left with \(27(a^6)(b^4)+12(a^2)(b^3) \iff \dfrac{27}{3}(a^6)(b^4)+\dfrac{12}{3} (a^2)(b^3) \\3(9a^6b^4+4a^2b^2) \)

OpenStudy (anonymous):

factor this: \(2y+8y^3\)

OpenStudy (anonymous):

2y(6+y) ??

OpenStudy (anonymous):

close. \(2y(1+4y^2)\) I'll step through it now

OpenStudy (anonymous):

How...

OpenStudy (anonymous):

let's start with just the numbers 2 and 8, the factors of 2: 1 2 1 2 4 8 so what is the greatest common factor?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so now we divide each term by 2

OpenStudy (anonymous):

my bad i was getting a snack.

OpenStudy (anonymous):

share :(

OpenStudy (anonymous):

\(2y+8y^3\implies \dfrac{2y}{2}+\dfrac{8y^3}{2}\\2(y+4y^3)\) there's another step we can do here

OpenStudy (anonymous):

~gives half of my chicken cordon blu~

OpenStudy (anonymous):

I do not under stand how to do that..

OpenStudy (anonymous):

what does \(y^2\) mean?

OpenStudy (anonymous):

y*y

OpenStudy (anonymous):

and so we have 2 terms here a) y b) y*y we have a common term of y that we can pull out

OpenStudy (anonymous):

you jusy y....

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

I Am 50%*2 confused...

OpenStudy (anonymous):

\(2y(\dfrac{y}{y}+\dfrac{4y^3}{y})\) in order to pull out a y, we have to divide each term by y. and so now we simplify the stuff int he parentheses and we get \(2y(1+4y)

OpenStudy (anonymous):

\(2y(1+4y)\)

OpenStudy (anonymous):

So, y/y= 1 and if we cancel out the ys on the other we get... 4 y... so it would be 2y(1+4y)

OpenStudy (anonymous):

and so that's the basic process we do for your problem.